OpenStudy (cwtan):

\[\int{cos^2 x}\]=? and \[\int{sin^3 x}\]=? Explain the steps you have done.

OpenStudy (mimi_x3):

\[\int\limits\frac{1+\cos2x}{2} dx\]

OpenStudy (mimi_x3):

Since, \[\cos^2x = \frac{1+\cos2x}{2} \]

OpenStudy (mimi_x3):

\[\sin^2x = \frac{1-\cos2x}{2} \]

OpenStudy (cwtan):

it's \[\Huge\sin^3 x\]

OpenStudy (mimi_x3):

\[\int\limits \sin^{3} xdx => \int\limits \sin^{2} xsinxdx => \int\limits(1-\cos^{2} x)(sinxdx\]

OpenStudy (cwtan):

Thank you I get it

OpenStudy (mimi_x3):

np